OpenStudy (anonymous):

A $7000 investment earns $840 annual simple interest in one year. What is the annual interest rate?

OpenStudy (texaschic101):

Interest = Principle * rate * time I = PRT, but since we are looking for rate (R) let me rearrange this formula. R = I / PT P = 7000 I = 840 T = 1 now we sub... R = 840/(7000 * 1) R = 840/7000 R = 0.12 but they want it as a percent so multiply it by 100 R = 12%

OpenStudy (anonymous):

Thank You for your help.:)

OpenStudy (texaschic101):

anytime :)